OpenStudy (anonymous):

Working together, two pumps can drain a certain pool in 3 hours. If it takes the older pump 7 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own? Do not do any rounding.

OpenStudy (anonymous):

I guess we assume they both drain at a constant rate.

OpenStudy (anonymous):

4 I think?

OpenStudy (anonymous):

assuming yes :)

OpenStudy (***[isuru]***):

hi, let's say that the volume of the pool is "x" liters then the rate at the older pump drain the pool is x/7 liters per hour and let's assume that the time newer pump will take to drain the pool "h" hours then the rate at the newer pump drain the pool is x/h liters per hour If they start working together then the rate at which they drain the pool is x/7 + x/h = (hx + 7x)/7h = x(7+h)/7h now we know the time take when 2 pumps work together... so.. rate x time = volume of the pool x(7+h)/7h * 3 = x 3 (7 + h) = 7h 21 = 4h h = 21/4 hours yep!! that's the answer!! hope this will help ya!!!

OpenStudy (anonymous):

We can let \(V\) be the volume of the pool. \[ 3(d_1+d_2)=V = 7(d_1) \]

OpenStudy (anonymous):

So \[ 3d_1+3d_2=7d_1 \]

OpenStudy (anonymous):

thanks team

OpenStudy (***[isuru]***):

u r welcome!!